# Effective Interactions in Quasi-One-Dimensional Dipolar Quantum Gases

### v1

@@ III.1 Other system geometries
We have by now analyzed the impact of the dipolar potential for the specific case of dipoles oriented parallel to the trap axis, where the interaction is predominantly attractive. A natural question arises as to whether corrections persist when the dipole moments are oriented at other angles, including the “magic angle”, where the dipole-dipole interaction is expected to vanish in the 1D model as predicted by Eq. (10). To address this scenario, we studied the adiabatic potentials in the harmonic oscillator basis, which are presented in Fig. 5A for the magic case and  5B for perpendicular dipole orientation, which leads to asymptotically repulsive interaction studied before by Guan et al. (2014). Note that now the trap axis is not aligned with the dipole, leading to couplings between channels with different $m$. For identical bosons, we are interested in symmetric wavefunctions, which can be realized with both $m$ parities. Based on previous results, it is now not surprising that the lowest adiabatic potential differs from zero even for the magic angle case. In fact, it has a range comparable to the trap lengthscale and decays at large distance approximately as $x^{-6}$. While this is a short-range potential with a rather weak amplitude, it can be expected to shift $a_{1D}$ from the value predicted by Eq. (12). This shift should be particularly pronounced for weak effective interactions. For the repulsive dipole configuration, the lowest adiabatic potential also becomes attractive at short range, shifting the confinement-induced resonance position.
Figure 5: Analysis of the 1D system with dipole-dipole interaction (A) at the “magic angle” $\theta_{F}=\arccos(1/\sqrt{3})\approx 55\degree$
. The adiabatic potentials diagonalized in the basis of harmonic oscillator modes, (B) with dipole orientation perpendicular to the trap axis.
@@ IV Discussion and Conclusions
We suggest using the developed framework as a reliable foundation for extending the analysis to molecular dipolar systems Bigagli et al. (2024). Correct treatment of dipole-dipole interactions might be crucial in theoretical analysis of quantum phases highly sensitive to the scattering parameters such as the quantum droplets Schmidt et al. (2022); Kopyciński et al. (2023); Langen et al. (2025); Łebek et al. (2025).

### v2

@@ III.1 Other system geometries
We have by now analyzed the impact of the dipolar potential for the specific case of dipoles oriented parallel to the trap axis, where the interaction is predominantly attractive. A natural question arises as to whether corrections persist when the dipole moments are oriented at other angles, including the “magic angle”, where the dipole-dipole interaction is expected to vanish in the 1D model as predicted by Eq. (10). To address this scenario, we studied the adiabatic potentials in the harmonic oscillator basis, which are presented in Fig. 5A for the magic case and  5B for perpendicular dipole orientation, which leads to asymptotically repulsive interaction studied before by [67]. Note that now the trap axis is not aligned with the dipole, leading to couplings between channels with different $m$. For identical bosons, we are interested in symmetric wavefunctions, which can be realized with both $m$ parities. Based on previous results, it is now not surprising that the lowest adiabatic potential differs from zero even for the magic angle case. In fact, it has a range comparable to the trap lengthscale and decays at large distance approximately as $x^{-6}$. While this is a short-range potential with a rather weak amplitude, it can be expected to shift $a_{1D}$ from the value predicted by Eq. (13). This shift should be particularly pronounced for weak effective interactions. For the repulsive dipole configuration, the lowest adiabatic potential also becomes attractive at short range, shifting the confinement-induced resonance position.
Figure 5: Analysis of the 1D system with dipole-dipole interaction (upper) at the “magic angle” $\theta_{F}\approx 55\degree$. The adiabatic potentials diagonalized in the basis of harmonic oscillator modes, (lower) with dipole orientation perpendicular to the trap axis.
@@ IV Discussion and Conclusions
We suggest using the developed framework as a reliable foundation for extending the analysis to molecular dipolar systems [77]. Correct treatment of dipole-dipole interactions might be crucial in theoretical analysis of quantum phases highly sensitive to the scattering parameters such as the quantum droplets [78, 79, 80, 81].